# theorem]Definition

### v1

@@ 3.3 OTPrune Selector
Table 1: Comparison of OTPrune (ours) with one-shot, adaptive, and finetuning methods on 11 benchmarks. DivPrune matches the PruMerge selection ratio. Following DivPrune, one-shot methods use a fixed token budget (9.8% retained), with TFLOPs acting as a complementary hardware-agnostic reference. The Avg. Rank $\downarrow$ is computed across all methods, and the One-shot Rank $\downarrow$ only among one-shot methods. OTPrune consistently achieves the best ranks. $\bullet$}: Finetuning used, $\bigtriangleup$}: Calibration used.
@@ 5.1 Experimental Settings
Datasets. We selected a comprehensive set of common tasks and 11 datasets aimed at multimodal reasoning and understanding involving ScienceQA-IMG (SQA) [32], POPE [24], MME [57], MMB [31], GQA [18], MMMU [53], Flicker30k [37], SeedBench (SEEDB) [22], Nocaps [1], OKVQA [33], and COCO-2017 [26]. These datasets encompass a wide range of tasks, including captioning, multiple-choice Question Answering (QA), and open-ended QA based on text and image inputs.
Evaluation metrics. Consistent with prior works [4], CIDEr score [43] is used for evaluating captioning tasks, and Exact Match (EM), Accuracy (Acc), F1, Perception Score (P-score) [57] and GPT-assisted [13] score are used for QA tasks. Furthermore, Wu-Palmer similarity (WUPS) score [48] and GPT-assisted score [13] is used for open-ended QA. For all task performance metrics used in this paper, higher values indicate better performance. Following the earlier works in  [10, 50, 28], we also report the computational requirement, measured in TFLOPs (consistency with [4]). Lower values indicate better results.
Baselines and Models. Following [4], we consider five baselines: FastV [10], PruMerge [40], VTW [28], FitPrune [50], $\text{M}^{3}$ [6], and DivPrune [4]. Among these, we consider FastV, VTW, DivPrune, and PruMerge as our main competitors as they are plug-and-play for each sample and do not rely on any further costly finetuning or calibration process. FastV, VTW, and DivPrune are one-shot, meaning they require a predefined pruning ratio shared across all samples, whereas PruMerge is adaptive, dynamically determining how many and which tokens to retain based on each image’s content rather than relying on a fixed threshold or ratio. However, for the sake of completeness, we also report performance comparison with respect to one finetuning-based ($\text{M}^{3}$) and one calibration-based (FitPrune) method. Note that, VTW, by default, requires calibration to determine the best layer for a given task. However, doing that does not allow us to set a specific TFLOP ratio, complicating the comparison. Hence, whenever required, we disable the calibration of VTW to select the layer that matches the FLOP requirement of a particular experiment.
Implementation. We test OTPrune and the baselines with different popular LMMs, involving LLaVA 1.5-7B [30] [^2] https://huggingface.co/liuhaotian/llava-v1.5-7B, LLaVA 1.5-13B [30] [^3] https://huggingface.co/liuhaotian/llava-v1.5-13b, and LLaVA 1.6-7B [^4] https://huggingface.co/liuhaotian/llava-v1.6-vicuna-7b (also known as LLaVA-NeXT [29]), to demonstrate the generality of DivPrune. For each tested model and task, we report only the relevant subset of baseline that is applicable to that specific model and task, alongside our results. All the tested LMMs used CLIP vision encoder [39]. LLaVA 1.5 model uses 576 visual tokens to represent images. LLaVA 1.6 converts each image into a varying number of patches, resulting in 3-5 times more visual tokens compared to LLaVA 1.5. We used 1 $\times$ A100 GPU with 80GB VRAM for all the experiments in this paper. Additionally, we used the lmms-evals package [55] for running these benchmarks for all the baselines and models. All results are obtained with a batch size of 1. For the metrics that require ChatGPT API access, the model is set to “gpt-4o-mini”.

### v2

@@ 3.3 OTPrune Selector
Table 1: Comparison of OTPrune with diverse baselines across 11 multimodal benchmarks. DivPrune aligns its token retention rate with PruMerge. For fair evaluation, all one-shot methods strictly retain 9.8% of visual tokens, using TFLOPs to indicate computational efficiency. The Avg. Rank $\downarrow$ assesses all evaluated methods, while the One-shot Rank $\downarrow$ strictly compares one-shot techniques. Our OTPrune consistently delivers the best performance. ($\bullet$}: requires finetuning; $\bigtriangleup$}: requires calibration).
@@ 5.1 Experimental Settings
Datasets. To rigorously evaluate our framework, we employ 11 distinct datasets covering a broad spectrum of multimodal reasoning and understanding benchmarks: ScienceQA-IMG (SQA) [32], POPE [24], MME [57], MMB [31], GQA [18], MMMU [53], Flicker30k [37], SeedBench (SEEDB) [22], Nocaps [1], OKVQA [33], and COCO-2017 [26]. This diverse collection allows us to assess capabilities across varied domains, ranging from image captioning to both multiple-choice and open-ended visual question answering scenarios.
Evaluation metrics. Following established evaluation protocols [4], we measure image captioning quality using the CIDEr metric [43]. For question-answering paradigms, model outputs are quantified via Exact Match (EM), Accuracy (Acc), F1, Perception Score (P-score) [57], and GPT-assisted evaluations [13]. Additionally, open-ended QA responses are judged utilizing the Wu-Palmer similarity (WUPS) [48] alongside GPT-assisted scoring [13]. Throughout our analysis, ascending values in these metrics consistently reflect superior predictive capability. To gauge efficiency, similar to previous studies [10, 50, 28], computational cost is recorded in TFLOPs (aligning with [4]), where lower numbers designate a more efficient process.
Baselines and Models. We compare OTPrune against several established token pruning techniques [4], specifically FastV [10], PruMerge [40], VTW [28], FitPrune [50], $\text{M}^{3}$ [6], and DivPrune [4]. Our primary analysis focuses on FastV, VTW, DivPrune, and PruMerge as our main competitors, as these offer plug-and-play functionality without the overhead of subsequent fine-tuning or dataset calibration. Within this group, FastV, VTW, and DivPrune operate under a one-shot paradigm (utilizing a globally shared pruning ratio), while PruMerge functions as an adaptive method (determining token retention dynamically per image). To ensure a comprehensive evaluation, we also provide comparative metrics against FitPrune (a calibration-dependent approach) and $\text{M}^{3}$ (a fine-tuning-based framework). It is important to highlight that VTW intrinsically relies on a calibration step to pinpoint the optimal pruning layer. Since this standard behavior prevents alignment with a fixed TFLOP target, we intentionally deactivate VTW’s calibration phase in our tests. Instead, we manually select a layer that forces VTW to satisfy the specific FLOP constraints of the given experiment, ensuring a fair, resource-matched comparison.
Implementation. To illustrate the broad applicability and generality of OTPrune, we deploy our algorithm and competing baselines across multiple widely used LMM architectures, including LLaVA 1.5-7B [30] [^2] https://huggingface.co/liuhaotian/llava-v1.5-7B, LLaVA 1.5-13B [30] [^3] https://huggingface.co/liuhaotian/llava-v1.5-13b, and LLaVA 1.6-7B [^4] https://huggingface.co/liuhaotian/llava-v1.6-vicuna-7b (alternatively referred to as LLaVA-NeXT [29]). In our tabulated results, we only display the baseline comparisons that are directly compatible with the specific model and task combination being evaluated. Architecturally, every LMM in our study incorporates the CLIP vision encoder [39]. While the LLaVA 1.5 variant maps images into a fixed grid of 576 tokens, LLaVA 1.6 scales the patch extraction dynamically, yielding a token sequence approximately 3 to 5 times longer than its predecessor. Hardware-wise, all benchmarks were executed on a single 1 $\times$ A100 GPU with 80GB VRAM. The inference and evaluation pipelines were constructed using the lmms-evals package [55], with tests running uniformly at a batch size of 1. Finally, for automated assessments dependent on the ChatGPT API, queries were processed via the “gpt-4o-mini” engine.